OpenStudy (anonymous):

A cylinder and a rectangular prism have the same volume and same height. The base of the prism is a square with side length 8 cm. What is the approximate radius of the cylinder? A. 2.5 cm B. 4.0 cm C. 4.5 cm D. 20.4 cm

OpenStudy (anonymous):

help

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

thx

OpenStudy (jdoe0001):

do you know what a rectangular prism is?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

draw a picture for starters.

OpenStudy (anonymous):

how

OpenStudy (anonymous):

im 3d and color blind

OpenStudy (anonymous):

or you can set up an equation: l*w*h=pie*r^2*h 8*8*h=pie*r^2*h

OpenStudy (anonymous):

64h=pie*r^2*h

OpenStudy (anonymous):

yes i think pi*r^2*h

OpenStudy (jdoe0001):

\(\bf \textit{volume of a cylinder}=\pi r^2 h \\ \quad \\ \textit{volume of a rectangular prism}=length\cdot width\cdot height\implies lwh \\ \quad \\ \textit{their volume is the same, thus }\pi r^2 h\iff lwh \\ \quad \\ \textit{base of the prism will be }length\cdot width=lw = 8\cdot 8=64 \\ \quad \\ \textit{thus we could say that }\pi r^2 h=lwh\implies \pi r^2 h=64h\) solve for "r"

OpenStudy (anonymous):

oh ok thx